Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
nardecchia, rocchi - 2:29 Wednesday 13 March 2019 (45242) Print this report
TCS CO2 benches check
Profiting of the activity on SDB2, we worked on the CO2 benches to check the general conditions of the CO2 beams.
For both benches we checked:
-the alignement and the superpostion of the DAS beams with the CH one (without the lens in order to have a small beam),
-the DASes beam shapes
-the ADC offsets and the power calibration.
Everything is okay and no changes have been observed wrt August 2018 when the DASes have been fine tuned.
As final check of the alignment of the DASes on the CPs we started the measurements with the HWS.
We began with the WI.
At around 21 UTC the ITF locked in recombined, we started the HWS-INJ acquisition at 21.51 UTC and we turned on the WI RH at 21.56 UTC to evaluate the center of the TM on the HWS-INJ map.
To do this, we blocked the SLED of HWS-DET on the bench by closing the shutter.
At 22.56 UTC the RH has been turned off. The TM centre was clearly visibile in the HWS-INJ map (see fig. 1).
At this poi we turned on the CH (still without lens) to evaluate the misalingnement of the DAS wrt the TM centre:
23.43 UTC: CH ON
At the beginning the CH appeared not perfectly centered in the HWS-INJ map, as shown in figure 2. We managed to recenter it. The starting misalignment was about 0.5 cm in both X and Y.
The final, current situation is shown in fig. 3.


Images attached to this report
Comments to this report:
nardecchia, rocchi - 17:36 Wednesday 13 March 2019 (45252) Print this report
This morning we checked the alignement of the DASes on the NI.
With the ITF in recombined, we started the acquisition of HWS-DET and we switched on the NI RH at 9.07 UTC to evaluate the center of the TM on the HWS-DET map.
To do this, we turned off the SLED of HWS-INJ and we opened the shutter on HWS-DET. At 10.17 UTC the RH has been turned off. The TM centre was clearly visibile in the HWS-DET map (see fig. 1).
At this point we turned on the CH (still without lens) to evaluate the misalingnement of the DAS wrt the TM centre:
11.02 UTC: CH ON.
At the beginning the CH appeared not centered in the HWS-DET map, as shown in figure 2. We managed to recenter it. The starting misalignment was about 1.5 cm in X and 0.3 cm in Y.
The final, current situation is shown in fig. 3.
Images attached to this comment
Search
×

Warning

×